OpenStudy (anonymous):

how you would factor 15x^4+ 50x^3– 40x^2?

OpenStudy (anonymous):

5x^2(3x^2 + 10x - 8) does this help ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Yo dawg, I herd u like factoring, so I put a greatest common factor in your factors so you can factor while you factor.

OpenStudy (anonymous):

lol XD

OpenStudy (anonymous):

so can you factor this : 5x^2(3x^2 + 10x - 8) completely now ?

OpenStudy (anonymous):

5x^2(x+4)(3x-2)?

OpenStudy (anonymous):

Now find the values of x for which y becomes zero...

OpenStudy (anonymous):

I mean *the equation becomes zero

OpenStudy (anonymous):

Did you understand what I said?

OpenStudy (anonymous):

not rly :/

OpenStudy (anonymous):

Let 5x^2(x+4)(3x-2) = y When does y become zero?

OpenStudy (anonymous):

i dunno

OpenStudy (anonymous):

y becomes zero when 5x^2=0 or x+4=0 or 3x-2=0

OpenStudy (anonymous):

So, y becomes zero when x=0 or x=-4 or x=2/3..These are the roots of the equation.. got it?

OpenStudy (anonymous):

sorta

OpenStudy (anonymous):

Ask now if you have any doubt...

OpenStudy (anonymous):

I think I got it.........I'm just thinking

OpenStudy (anonymous):

-24 (a*c = -24). This can be achieved with 12x-2x, as 12*-2 = -24. =5x^2(3x^2+12x-2x-8) = 5x^2[(3x(x+4)-2(x+4)] = 5x^2(3x-2)(x+4) Factors are 5x^2, (3x-2), (x+4)

OpenStudy (anonymous):

she just needed to factor it .. 5x^2(x+4)(3x-2) this is a right answer

OpenStudy (anonymous):

Oh sorry:( I think I messed it up.